in a recent election Corrie Broewn recieved 13,517 more votes than Bill Randall. if the total number of votes was 118,985, find the number of votes for each candidate
Answer by rfer(16322)

x+x+13517=118985
2x=118985-13517
2x=105468
x=52734
x+13517=66251